i need proper answer please

Consider a closed economy to which the Keynesian-cross analysis applies. Consumption is given by the equation C = 200 + 0.75(Y – T), where Y represents income and T represents taxes. Planned investment is 300, as are government spending and taxes.
a. What is the planned expenditure function? What is equilibrium level of income?
b. If Y is 1,500, what are planned aggregate expenditures? What is inventory accumulation or
deacumulation? Should equilibrium Y be higher or lower than 1,500?(equilibrium; hint; planned expenditure=Y, and inventory =Y-planned expenditure)
c. What are equilibrium consumption, private saving, public saving, and national saving?hint:private saving=Y-C-T
public saving=T-G
equilibrium saving=Y-C-T

d. How much does equilibrium income decrease when G is reduced to 200?
What is the multiplier for government spending?
1
Expert's answer
2018-04-19T09:41:08-0400
C =200 + 0.75(Y - T), I = G = T = 300.
a. what is the planned expenditure function? what is equilibrium level of income?

The planned expenditure function,
PE = C + I + G = 200 + 0.75(Y - T) + 300 + 300 = 800 + 0.75(Y - 300) = 800 + 0.75Y - 225 = 575 + 0.75Y
So, the planned expenditure function, PE = 575 + 0.75Y
If Y is 0, PE = 575. So, the vertical intercept (PE-intercept) of the expenditure function is 575.
The equilibrium condition:
Y = PE
or, Y = 575 + 0.75Y
or, Y - 0.75Y = 575
or, 0.25Y = 575
or, Y = 2300
So, equilibrium level of income is 2300.
b. if y is 1500, what are planned aggregate expenditures? what is inventory accumulation or deaccumulation? should equilibrium Y be higher or lower than 1500?(equilibrium; hint; planned expenditure=Y, and inventory =Y-planned expenditure)

If Y is 1500, then Planned spending,
PE = C + I + G = 200 + 0.75(Y - T) + 300 + 300 = 800 + 0.75(1500 - 300) = 800 + 0.75*1200 = 800 + 900 = 1700.
So, at Y = 1500, planned spending is 1700.
Since at Y = 1500, planned spending (1700) is higher than current output, there will be unplanned inventory decumulation. The amount of inventory decumulation = (PE - Y) = (1700 - 1500) = 200. At Y = 1500, firms meet the excess demand by supplying goods from their inventories, resulting into an unplanned reduction in their stock of goods. Firms would respond to this decline in inventories by producing more goods. So, the equilibrium Y must be higher than 1500.
c. what are equilibrium consumption, private saving, public saving, and national saving? hint: private saving = Y - C - T, public saving = T - G

The equilibrium consumption, C = 200 + 0.75(Y - T) = 200 + 0.75(2300 - 300) = 200 + 0.75*2000 = 1700
So, the equilibrium consumption is 1700.
The equilibrium private saving, SP = Y - C - T = 2300 - 1700 - 300 = 300
So, the equilibrium private saving is 300.
The equilibrium public saving, SG = T - G = 300 – 300 = 0
So, the equilibrium public saving is 0.
The equilibrium national saving, S = Y - C - T = 2300 - 1700 - 300 = 300
So, the equilibrium national saving is 300.

d. how much does equilibrium income decrease when G is reduced to 200? what is the multiplier for government spending?
The government spending multiplier = 1/(1 - MPC) = 1/(1 - 0.75) = 4.
So, the government spending multiplier is 4. We can use this multiplier to find the change in the equilibrium income.
Change in equilibrium income = Government spending multiplier * Change in G = 4*(-100) = -400.
So, the equilibrium income decreases by 400 when G is reduced to 200.